OpenStudy (anonymous):

which of the following is the equation of a line perpendicular to the line Y= -3/2x+4 and passes through the point (3,9) A) -2x-3y=-21 B) -2x+3y=21 C) 2x+3y=21 D) 2x-3y=21

OpenStudy (misty1212):

a) find the slope of \(y=-\frac{3}{2}x+4\) well you don't really find it, you just say it do you know it?

OpenStudy (misty1212):

if you do not, then say so and i will tell you

OpenStudy (anonymous):

-3/2

OpenStudy (misty1212):

ok good now the perpendicular line will have the slope that is the "negative reciprocal" what is that?

OpenStudy (anonymous):

3/2

OpenStudy (misty1212):

you didn't flip it

OpenStudy (anonymous):

2/3

OpenStudy (misty1212):

ok good now the point slope formula gives \[y-9=\frac{2}{3}(x-3)\]

OpenStudy (anonymous):

so what do i do next? im confused :(

OpenStudy (misty1212):

write it in standard form first multipy all by 3 to get \[3y-27=2(x-3)\] then expand on the left

OpenStudy (misty1212):

\[3y-27=2x-6\] then put it all on one side of the equal sign

OpenStudy (misty1212):

\[2x-3y=-21\] which is not an option so maybe \[-2x+3y=21\]

OpenStudy (misty1212):

oh yeah that is a choice